# Fernando Castor

### v1

@@ 3.3 Collecting energy data
Mac systems provide a command-line utility that reports on CPU and GPU power usage called powermetrics. It is able to collect data at a sampling rate of up to 1kHz (one observation per milisecond). When Macs used Intel processors, powermetrics would leverage Intel’s RAPL interface [24] to get fairly good estimates of processor and RAM energy use. After Apple transitioned into using its own proprietary processor on Macs, the details of how powermetrics collects data became less clear, as Apple does not disclose them and the man page [^12] https://www.unix.com/man-page/osx/1/powermetrics/ is outdated (as of July 16, 2024). Notwithstanding, considering that Apple has total control over the design of its processors and operating system and many years of experience in building mobile devices, we assume that the results reported by it are at least reliable enough to be used to optimize benchmarks and applications for energy efficiency, though maybe not to compare different machines. Before we perform the experiment, we run powermetrics for a few seconds to gauge the idle power usage of the device. More specifically, we run the following command on the Mac terminal:
@@ 4.3 Analytical models
Modeling is a complex topic and we have no hope of even scratching the surface in this document. Models can vary widely in terms of complexity, granularity, mathematical framework, required data, moment (predictive, on-line, or a posterior estimation), and produced output. Here we focus on a specific type of modeling solution. The idea is to enable developers to compare alternative software solutions without directly considering the specific number of joules that each one requires under a certain workload. This type of solution has appeared multiple times in the literature [17, 29, 32] and is easy to use or even build from scratch. In addition, software developers are not used to thinking in terms of joules, with lack of appropriate KPIs being cited as a major obstacle to taking energy efficiency into account in software development [31]. On top of this, hardware platforms can impact energy consumption differently [21, 29]. Thus, placing the emphasis on the ability to compare results for software is a practical approach that circumvents these issues.
@@
This modeling approach can be applied in a non-obvious but simple manner to systems that employ deep learning and large language models. In this scenario, the “operation” of interest can be all floating point operations, which are usually executed by GPUs. We can estimate the number of floating point operations required to perform a certain task. The rate of floating operations per second (FLOPS) that contemporary GPUs can perform is publicly disclosed by their manufacturers [^22] https://www.nvidia.com/en-us/geforce/graphics-cards/40-series/rtx-4080-family/. GPUs are responsible for the bulk of the work to train and use deep learning systems. Since that work consists of performing a large number of floating point operations, measuring the time spent in these tasks and multiplying it by the number of FLOPS of the underlying GPU gives us an approximate measure that we can use to compare machine learning solutions, in terms of their energy footprint. If the execution environment is kept unchanged, this is a reasonable metric [12]. If factors such as the hardware configuration, deep learning framework or library [14, 36], runtime system [2], and machine learning task [2, 21, 14] can vary across executions, results may differ significantly.
@@ 5 Further study
The goal of this document is not to talk about tools to gather data about energy efficiency. Cruz and his collaborators [9] wrote a nice post about this topic. It is a bit outdated, e.g., it does not talk about (as of July 16, 2024) solutions such as the Snapdragon Profiler (which replaced the now-defunct Trepn) [^23] https://www.qualcomm.com/developer/software/snapdragon-profiler nor the On-Device Power Rails Monitor available in recent Android smartphones, and talks about discontinued solutions, such as Intel Power Gadget. Nevertheless, it is still a useful resource if the reader is interested in a hands-on perspective, written by folks who have experience in using these tools.
@@ 6 Acknowledgments
We would like to thank Ivano Malavolta for kindly reading and providing feedback on a preliminary version of this document. We would also like to thank Arne Tarara for writing very interesting articles on low level aspects of Green Software Developmentand, more specifically, for an enlightening discussion about SGX and energy filtering in Intel machines.

### v2

@@ 3.3 Collecting energy data
Mac systems provide a command-line utility that reports on CPU and GPU power usage called powermetrics. It is able to collect data at a sampling rate of up to 1kHz (one observation per milisecond). When Macs used Intel processors, powermetrics would leverage Intel’s RAPL interface [26] to get fairly good estimates of processor and RAM energy use. After Apple transitioned into using its own proprietary processor on Macs, the details of how powermetrics collects data became less clear, as Apple does not disclose them and the man page [^12] https://www.unix.com/man-page/osx/1/powermetrics/ is outdated (as of July 17, 2024). Notwithstanding, considering that Apple has total control over the design of its processors and operating system and many years of experience in building mobile devices, we assume that the results reported by it are at least reliable enough to be used to optimize benchmarks and applications for energy efficiency, though maybe not to compare different machines. Before we perform the experiment, we run powermetrics for a few seconds to gauge the idle power usage of the device. More specifically, we run the following command on the Mac terminal:
@@
A relevant limitation of the employed methodology is that it only considers dynamic power. Leakage power, the power that stems from current passing through a circuit when there should be none, can also be significant [13]. We are not aware of any previous paper in the area of Green Software Developmentthat explicitly accounts for leakage power.
@@ 4.3 Analytical models
Modeling is a complex topic and we have no hope of even scratching the surface in this document. Models can vary widely in terms of complexity and granularity. For example, a state-of-the-art analytical model capable of estimating GPU power, such as AccelWattch [25], can consider a wide range of factors, including “constant, static, and dynamic power consumption in the presence of DVFS, thread divergence, intra-warp functional unit overlap, variable SM [Streaming Multiprocessor] occupancy, and power gating.”. Conversely, to model the energy footprint of the execution of a GPU heavy application, we can simply multiply the rate of floating operations per second (FLOPS) of the underlying GPU by its execution time (more on this later). Other factors to consider when building a model for energy include the employed mathematical framework, required data, moment (predictive, on-line, or a posterior estimation), and produced output.
In this section we focus on a specific type of modeling solution. The idea is to enable developers to compare alternative software solutions without directly considering the specific number of joules that each one requires under a certain workload. This type of solution has appeared multiple times in the literature [18, 31, 34] and is easy to use or even build from scratch. In addition, software developers are not used to thinking in terms of joules, with lack of appropriate KPIs being cited as a major obstacle to taking energy efficiency into account in software development [33]. On top of this, hardware platforms can impact energy consumption differently [22, 31]. Thus, placing the emphasis on the ability to compare results for software is a practical approach that circumvents these issues.
@@
This modeling approach can be applied in a non-obvious but simple manner to systems that employ deep learning and large language models. In this scenario, the “operation” of interest can be all floating point operations, which are usually executed by GPUs. We can estimate the number of floating point operations required to perform a certain task. The rate of FLOPS that contemporary GPUs can perform is publicly disclosed by their manufacturers [^22] https://www.nvidia.com/en-us/geforce/graphics-cards/40-series/rtx-4080-family/. GPUs are responsible for the bulk of the work to train and use deep learning systems. Since that work consists of performing a large number of floating point operations, measuring the time spent in these tasks and multiplying it by the number of FLOPS of the underlying GPU gives us an approximate measure that we can use to compare machine learning solutions, in terms of their energy footprint. If the execution environment is kept unchanged, this is a reasonable metric [12]. If factors such as the hardware configuration, deep learning framework or library [15, 38], runtime system [2], and machine learning task [2, 22, 15] can vary across executions, results may differ significantly.
@@ 5 Further study
The goal of this document is not to talk about tools to gather data about energy efficiency. Cruz and his collaborators [9] wrote a nice post about this topic. It is a bit outdated, e.g., it does not talk about (as of July 17, 2024) solutions such as the Snapdragon Profiler (which replaced the now-defunct Trepn) [^23] https://www.qualcomm.com/developer/software/snapdragon-profiler nor the On-Device Power Rails Monitor available in recent Android smartphones, and talks about discontinued solutions, such as Intel Power Gadget. Nevertheless, it is still a useful resource if the reader is interested in a hands-on perspective, written by folks who have experience in using these tools.
@@ 6 Acknowledgments
We would like to thank Ivano Malavolta and Negar Alizadeh for kindly reading and providing feedback on a preliminary version of this document. We would also like to thank Arne Tarara for writing very interesting articles on low level aspects of Green Software Developmentand, more specifically, for an enlightening discussion about SGX and energy filtering in Intel machines.